OpenStudy (anonymous):

Solve the system. then classify the system as consistent and independent, consistent and dependent or inconsistent x+y=3 5x+3y=1 please help and can you explain it :)

OpenStudy (anonymous):

x+y=3 (*-3)= -3x-3y=-9 5x+3y=1 ___________ 2x=-8, x=-4 -4+y=3 y=7 consistent dependent happy to help :P

OpenStudy (anonymous):

Thanks so much can ou help me with this one too? 3X-4Y=-5 6X+4Y=14

OpenStudy (anonymous):

oops. consistent INdependent. my bad

OpenStudy (anonymous):

3x-4y=-5 3*1-4y=-5, -4y=-8, 4y=8,y=2, consistent independent 6x+4y=14 __________ 9x=9, x=1

OpenStudy (anonymous):

tHANKKS :)